# MindGYM: Enhancing Vision-Language Models via Synthetic Self-Challenging Questions

### v1

@@ 3.2 Challenging Multi-Hop QA Synthesis
3.2.1 Cognitive Combination Principles
Synthesis Templates for In-depth Thinking.
Recall that in Section 3.1, we generate $k$ single-hop seed questions and proactively inform $\pi$ that these questions will be used for future cognitive combinations. Here, we explicitly specify some cognitive combination relationships into the synthesis requirement $P_{task}$ for multi-hop QAs, and let $\pi$ to self-challenging by figuring out how to combine the generated $\{q^{1}_{i},...q^{k}_{i}\}_{i\in N}$ into more complex questions $\{Q_{i}\}_{i\in N}$ conditioned on aforementioned $P_{cog}$ and $P_{context}$, such that the results is logically right, cognitively considered and match the most suitable relationships among the given ones.
Building on the single-hop seed questions generated in Section 3.1, we advance to multi-hop question synthesis by explicitly defining cognitive combination relationships within the task specification $P_{\text{task}}$. During this phase, the model $\pi$ is prompted to synthesize more complex questions $\{Q_{i}\}_{i\in N}$ by logically combining the pre-generated seed questions $\{q^{j}_{i}\}_{i=1,...,N;j=1,...,k}$ under constraints from $P_{\text{cog}}$ (cognitive framework) and $P_{\text{context}}$ (factual reference). This process aims to make $\{Q_{i}\}_{i\in N}$ logically valid, cognitively grounded, and optimally match the most relevant relationships from our predefined taxonomy.
Textual Cognitive Combination Types.
For $P_{\text{context}}$ containing only textual content, we define three cognitive combination forms after initial seed generation:
  • Comparison: Requires contrasting two or more concepts or scenarios.
  • Bridging: Connects disparate concepts through intermediate reasoning steps.
  • Temporal: Involves analyzing events across chronological contexts.
Multimodal Cognitive Combination Types.
3.2.2 Cognitive Manifestation
Structured Cognitive Processes & Reasoning Units.
Beyond specifying cognitive combination relationships in the task requirement $P_{\text{task}}$, we explicitly require $\pi$ to articulate its intermediate thought process $T_{i}$ leading to the final question $Q_{i}$. This support $T_{i}$ links the seed questions $\{q^{j}_{i}\}_{i\in N}$ into a logically coherent reasoning pathway, explicitly connecting steps to ensure interpretability and contextual consistency.
To balance cognitive context relevance and inference efficiency, $\pi$ is prompted to jointly generate $Q_{i}$, $A_{i}$, and $T_{i}$ in one inference step. However, raw responses may include fragmented or overly lengthy content, so we further prompt $\pi$ to segment outputs into distinct components: formal question-answer pairs $(Q_{i},A_{i})$ and their supporting rationale $T_{i}$. This results in modular cognitive units of varying granularity, enabling more flexible refinement of $\pi$’s reasoning abilities in subsequent training (more details are in later Section 3.3).
Cognitive Information Processing Dynamics.
Note that the synthesis framework explicitly involves $\pi$ processing both novel and existing knowledge during inference. By reintroducing context elements from $P_{\text{context}}$, accessing seed questions $\{q^{j}_{i}\}_{i\in N}$, and addressing advanced combination requirements in $P_{\text{task}}$, the model is incentivized to: (1). Explore diverse reasoning patterns embedded in $P_{\text{cog}}$, (2). Re-examine context details with renewed emphasis, and (3). Construct multi-hop answers using enriched relational constraints. This process yields more focused, nuanced questions $Q_{i}$ while eliminating reasoning shortcuts, forcing $\pi$ to engage in deeper inferential analysis.
3.3 Thinking-Induced Curriculum Fine-Tuning
Training Objectives.
Given the self-synthesized dataset containing triples $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$, we propose a “MindGym”-style training flow to progressively integrate cognitive guidance into the model $\pi$ through mixed masking strategies. This approach incrementally challenges $\pi$ by varying the exposure to question-answer pairs ($Q_{i}$, $A_{i}$), their supporting rationales ($T_{i}$), and masked input-output configurations.
For recap, $Q_{i}$ represents the complex multi-hop question, $A_{i}$ denotes the reference answer with structured reasoning, and $T_{i}$ records the detailed step-by-step cognitive trajectory. We construct four training datasets with distinct masking objectives:
$\mathcal{D}^{1}_{\text{train}}=\{(Q_{i},T_{i}\rightarrow A_{i})\}$ employs the masked input structure $\langle Q_{i},T_{i},[\mathbf{\text{MASK}}]\rangle$. The model predicts $A_{i}$ using $Q_{i}$ and explicit thinking guidance $T_{i}$, and the loss is defined as:
$\mathcal{L}_{1}=-\mathbb{E}_{(Q,T,A)\sim\mathcal{D}^{1}_{\text{train}}}\log\pi_{\theta}(A|Q,T).$ (1)
$\mathcal{D}^{2}_{\text{train}}=\{(Q_{i},A_{i}\rightarrow T_{i})\}$ uses the structure $\langle Q_{i},[\mathbf{\text{MASK}}],A_{i}\rangle$. The objective is to recover the hidden thinking path $T_{i}$, with loss:
$\mathcal{L}_{2}=-\mathbb{E}_{(Q,A,T)\sim\mathcal{D}^{2}_{\text{train}}}\log\pi_{\theta}(T|Q,A).$ (2)
$\mathcal{D}^{3}_{\text{train}}=\{Q_{i}\rightarrow(A_{i},T_{i})\}$ follows $\langle Q_{i},[\mathbf{\text{MASK}}]\rangle$, requiring simultaneous generation of $A_{i}$ and $T_{i}$:
$\mathcal{L}_{3}=-\mathbb{E}_{(Q,A,T)\sim\mathcal{D}^{3}_{\text{train}}}\log\pi_{\theta}(A,T|Q).$ (3)
$\mathcal{D}^{4}_{\text{train}}=\{Q_{i}\rightarrow A_{i}\}$ uses the same masking strategy as $\mathcal{D}_{train}^{3}$, with input $\langle Q_{i},[\mathbf{\text{MASK}}]\rangle$. This mimics standard supervised fine-tuning practice:
$\mathcal{L}_{4}=-\mathbb{E}_{(Q,A)\sim\mathcal{D}^{4}_{\text{train}}}\log\pi_{\theta}(A|Q).$ (4)
The curriculum progresses through these datasets sequentially: $\mathcal{D}^{1}\rightarrow\mathcal{D}^{2}\rightarrow\mathcal{D}^{3}\rightarrow\mathcal{D}^{4}$. This mirrors a guided learning pathway where $\pi$ first reproduces answers with external cognitive scaffolding ($\mathcal{D}^{1}$), then deduces thinking paths without full guidance ($\mathcal{D}^{2}$), gradually internalizes end-to-end reasoning ($\mathcal{D}^{3}$), and finally masters standalone Q-to-A mappings ($\mathcal{D}^{4}$). Empirical validation in Sec. 4.3 shows this strategy outperforms non-curriculum mixed training.
Discussion on Efficiency and Cost.
The MindGYM framework enhances $\pi$’s reasoning capabilities through $N\times 4$ synthesis inferences across stages ($P_{\text{context}}$, seed questions, $Q_{i}$, and $\mathcal{D}_{\text{train}}$). Despite this incremental setup, inference efficiency is maintained through amortized reasoning costs during training. A single inference suffices to achieve high model performance, as shown empirically in Section 4.3. Unlike test-time scaling approaches [38, 18], we move cognitive enhancement costs to the training phase. Compared to data-intensive methods (e.g., reinforcement learning) requiring extensive trial-and-error [19], our method maximizes cognitive depth leveraging structured synthetic data without resource-heavy exploration phases.
Datasets Multimodal Eval Text-Only Eval Avg
MMStar MathVista MathVision MM-Avg GSM8K MATH GPQA Text-Avg
Baslines Raw 64.00 69.30 24.67 52.66 83.62 67.60 31.83 61.02 56.84
LIMO (817) [43] 63.93 70.20 25.9 53.37 84.08 67.80 30.58 60.82 57.10
Open-O1 (77k) [35] 58.93 61.40 17.43 45.92 77.94 59.00 28.19 55.04 50.48
Open-O1 (400) [35] 63.87 69.70 23.66 53.08 84.31 69.00 29.7 61.00 57.04
Text MindGYM (400) 64.33 70.30 28.62 54.42 84.08 68.40 31.33 61.27 57.84
Image MindGYM-OKVQA (400) 63.60 69.20 28.29 53.70 84.00 68.20 31.08 61.09 57.40
MindGYM-SciQA (400) 64.20 69.80 25.33 53.11 83.93 69.00 31.45 61.46 57.29
@@ 3.3 Thinking-Induced Curriculum Fine-Tuning
Table 1: Performance of Qwen2.5-VL-7B on various evaluation benchmarks. The number following the dataset name indicates its size in terms of the number of samples used for training.
@@ 4 Experiments
To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative text and multimodal evaluation sets, comparing it with several competitive data synthesis methods and datasets. Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix A.
@@ 4.1 Experimental Settings
We conduct multi-modality data synthesis across textual and image contexts. For textual synthesis, data generation occurs exclusively through self-contained knowledge extraction from the base model, eliminating external data dependencies. Multimodal synthesis utilizes established scientific and reasoning corpora as visual primitives, with visual grounding from ScienceQA [22] and OK-VQA [25].
@@
To systematically assess MindGYM on VLMs across reasoning modalities, we incorporate:
  • Multimodal evaluation sets include MMStar [9], MathVista-Mini [23], and MathVision-Mini [37], designed to comprehensively assess the reasoning abilities in the multimodal scenario.
  • Text-based evaluation sets comprise GSM8K [10], MATH [15], and GPQA [31], focused on evaluating reasoning and mathematical capabilities.
The MM-Avg and Text-Avg metrics serve as unified criteria for assessing multimodal and text-based reasoning capabilities respectively, whereas Avg quantifies comprehensive performance spanning both unimodal and multimodal reasoning domains.
Models & Implementation.
The data synthesis of MindGYM leverages VLMs’ inherent long-context understanding and reasoning capabilities. We implement Qwen2.5-VL-7B-Instruct [4] as our base VLM since alternative models (e.g., InternVL2.5-7B) exhibit inferior generation quality under identical configurations (detailed in Appendix D). To maintain linguistic consistency with Qwen2.5-VL-7B-Instruct and consider the information density of reasoning data, we utilize Chinese for data synthesis. All experiments follow consistent protocols: during text-only training, only LLM-layers update parameters, while image-text joint training additionally activates Aligner-layers, keeping ViT-layers permanently frozen.
@@
As MindGYM generates self-contained adversarial training data without external sources, direct comparison with conventional data synthesis methods proves infeasible. We therefore select SOTA reasoning-oriented dataset works from two categories: (1) Multimodal reasoning using MMEvol’s [24] core subsets ScienceQA and DVQA, and (2) Text-based reasoning with LIMO [43], 817 curated logical chains, and Open-O1 [35], an SFT dataset for CoT activation as fundamental text reasoning baselines.
@@ 4.2 Main Results
Universal Superiority.
MindGYM demonstrates substantial performance advantages through comprehensive average metric evaluation in MindGYM-Text with a 54.42 MM-Avg-score and 57.84 Avg-score, and notably achieving 28.62 on MathVision with a 16% improvement over the Raw model. The proposed method significantly outperforms all baselines where even the strongest competitor LIMO trails by 1.05 points in MM-Avg. An interesting phenomenon emerges among baselines that Open-O1 (77K) suffers severe degradation with 6.74-point MM-Avg decline and 29.3% MathVision collapse despite employing a large number of training samples, exposing the vulnerability of conventional scaling approaches.
Cross-modal Transfer.
Generalizability & Robustness.
@@
MindGYM actives model potential using only 400 synthesized samples, achieving a 28.62 MathVision score that surpasses baselines trained with 10$\times$-100$\times$ more data. This contrasts sharply with Open-O1’s performance collapse despite massive data scaling, indicating that MindGYM significantly enhances the model’s reasoning capabilities while also ensuring that the model’s performance is not detrimentally affected.
These findings effectively validate that our approach addresses the issues of modality inconsistency, generalization, and computational efficiency mentioned in Section 2.

### v2

@@ 3.3 Challenging Multi-Hop QA Synthesis
Stage 3: Cognitive Composition with Adaptive Types.
Given the seed questions $\{q^{j}_{i}\}_{j=1}^{k}$ from Section 3.2, we treat them as updated context $P_{\text{context}}^{\text{single}}$ for this phase. The model $\pi$ is now prompted with the triplet $(P_{\text{cog}},P_{\text{context}}^{\text{single}},P_{\text{task}}^{\text{multi-hop}})$ to synthesize a more complex, cognitively grounded question $Q_{i}$. The task instruction $P_{\text{task}}^{\text{multi-hop}}$ specifies an abstract reasoning requirement (e.g., comparing entities, bridging concepts, or analyzing temporal relations), which guides the model to self-compose from prior seeds. Details of composition types are provided in Appendix B.
Stage 4: Structured Extraction.
To support modular training and interpretation, we encourage $\pi$ to format its output in three clearly separated blocks: the final multi-hop question $Q_{i}$, its answer $A_{i}$, and the structured rationale $T_{i}$. These thinking traces serve as valuable training signals for reasoning-intensive QA tasks, which is used for training in Section 3.4).
3.4 Usage and Discussion
Training with the Dataset.
In this section, we explore how to effectively utilize our cognitively annotated dataset $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$ by organizing the training process into a gradually evolving pathway that transitions from explicit external guidance ($T_{i}$) to self-contained reasoning. The process begins with guided answering, where the model receives both the question and the reasoning trace ($Q_{i}$, $T_{i}\rightarrow A_{i}$) to encourage answer generation aligned with structured thought. It then advances to reason reconstruction, where the model is given the question and answer and learns to infer the underlying rationale ($Q_{i}$, $A_{i}\rightarrow T_{i}$), reinforcing logical coherence. Next comes paired reasoning, which requires the model to jointly generate both the answer and rationale from the question alone ($Q_{i}\rightarrow A_{i}$, $T_{i}$), simulating independent reasoning under minimal supervision. Finally, in the autonomous solving phase, the model learns to directly produce the answer from the question ($Q_{i}\rightarrow A_{i}$), fully internalizing the reasoning process. This learning pathway reflects the natural progression of human cognitive development—starting from external support and moving toward self-directed problem solving. As demonstrated in Section 4.4, this structured progression leads to significantly better performance compared to training with uniform supervision.
Extensibility of the Dataset.
In addition to our primary focus on cognitively annotated textual data, we conduct preliminary experiments on extending our synthesis framework to the multimodal setting. As described in Appendix C, we leverage VLMs to generate multi-hop QA samples grounded in visual context. Specifically, we use OK-VQA okvqa and ScienceQA lu2022learn as anchor image sources to guide the synthesis process. The generated results summarized in Appendix C.3 show encouraging quality and reasoning depth, suggesting the potential of our framework beyond the text-only domain. Nonetheless, multimodal synthesis with VLMs remains an open challenge. Current reliance on static image datasets as anchors limits the diversity and scalability of visual reasoning. Looking forward, we plan to (1) incorporate richer and more diverse image datasets to better stimulate cognitive engagement, and (2) further explore fully generative pipelines where VLMs synthesize both the image and the associated reasoning task, enabling models to improve through self-generated multimodal experiences—a key step toward the self-evolution of models.
@@ 3.4 Usage and Discussion
Table 1: Performance of different models on various evaluation benchmarks. The number following the dataset name indicates its size in terms of the number of samples used for training.
Text Eval Multimodal Eval
Models Dataset (# of Samples) GSM8K Math GPQA MMStar MathVista MathVision Overall
[…]
Openo1-sft (400) openo1 84.31 69.00 29.70 63.87 69.70 23.66 57.04 (+0.20)
Openo1-sft (4k) openo1 77.94 59.00 28.19 58.93 61.40 17.43 50.48 (-6.36)
[…]
MMEvol-SciQA (106) luo2024mmevol 83.40 65.60 31.83 63.93 69.50 23.36 56.27 (-0.57)
MMEvol-DvQA (4k) luo2024mmevol 83.85 65.80 33.08 62.93 67.40 24.01 56.18 (-0.66)
[…]
MMEvol-SciQA (106) luo2024mmevol 95.60 81.00 42.42 69.00 73.80 36.84 66.44 (-1.14)
[…]
raw 86.50 76.30 33.84 69.00 73.20 33.22 62.46
Openo1-sft (400) openo1 89.39 77.20 43.43 69.00 73.60 32.57 64.20 (+1.74)
[…]
MMEvol-DvQA (4k) luo2024mmevol 88.55 74.60 40.40 68.47 71.60 32.24 62.65 (+0.19)
[…]
raw 89.16 72.60 47.47 72.40 72.80 36.51 65.16
[…]
Openo1-sft (4k) openo1 89.46 79.20 44.95 72.47 72.90 35.53 65.75 (+0.59)
LIMO (817) ye2025limo 89.16 72.60 47.47 72.40 72.80 36.51 65.16 (+0.00)
[…]
MMEvol-DvQA (4k) luo2024mmevol 89.92 74.00 46.46 70.93 71.70 30.92 63.99 (-1.17)
[…]
@@ 4 Experiments
To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative evaluation sets. Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix A.
@@ 4.1 Experimental Settings
Models & Implementation.
The data synthesis in MindGYM leverages VLMs’ inherent capabilities for long-context understanding and reasoning. We adopted four vision language models to verify the generality of our data synthesis method. These models are selected to cover both scaling within a single model family (Qwen2.5-VL 7B and 32B) and variation across different models (Qwen and InternVL series). For all experiments in the main text, we focus on text-only data synthesis, where only the LLM-layers are updated during training, while all vision-related layers remain frozen.
@@
We focus exclusively on synthetic textual data generation in the main body of our work, relying solely on self-contained knowledge extraction from the base model without external data sources. To verify the robustness and generality of our synthesis method, we employ the aforementioned models to perform self-consistent data generation. Each model independently synthesizes 400 Chinese samples. The choice of Chinese is driven by its linguistic complexity and higher information density, making the synthesis task more challenging.
@@
To systematically assess MindGYM on VLMs across reasoning modalities, we incorporate: (1) Text-based evaluation sets comprise GSM8K gsm8k , MATH math , and GPQA gpqa ; (2) Multimodal evaluation sets include MMStar MMStar , MathVista-Mini mathvista , and MathVision-Mini mathvision . The Overall quantifies comprehensive performance spanning both text and multimodal domains.
@@
As MindGYM generates self-contained adversarial training data without external sources, direct comparison with conventional data synthesis methods proves infeasible. We therefore select SOTA reasoning-oriented dataset works from two categories: (1) Text-based reasoning with LIMO ye2025limo , 817 curated logical chains, and Open-O1 openo1 , an SFT dataset for CoT activation as fundamental text reasoning baselines, and (2) Multimodal reasoning using MMEvol’s luo2024mmevol core subsets ScienceQA and DVQA. All baseline datasets are used in accordance with their original supervised fine-tuning setting. For fair comparison, we apply identical training protocols and hyperparameter configurations across all methods.
@@ 4.2 Main Results
Superiority across Different Tasks.
MindGYM demonstrates substantial performance advantages through comprehensive average metric evaluation. Notably, on the Qwen2.5-VL-7B, MindGYM achieves the Overall of 57.84, outperforming all competing baselines. It achieves the highest MathVision score (28.62), exceeding the Raw model by 16% and the strongest baseline LIMO by 2.72 points. An interesting phenomenon emerges among baselines that Open-O1 (4K) suffers severe degradation with 29.3% MathVision collapse despite employing a large number of training samples, exposing the vulnerability of conventional scaling approaches.
Superiority on Different Models.
MindGYM consistently demonstrates strong robustness across both diverse models and varying parameter scales. On the Qwen2.5-VL series, MindGYM improves the Overall performance from 57.84 to 64.33 on the 7B model, a relative gain of +0.74%–7.36%, and from 68.06 to 68.06 on the 32B model, reflecting a more modest but stable +0.48%–3.31% improvement. Similarly, for InternVL3, the Overall score rises from 64.95 to 64.95 on the 14B model (+0.12%–2.30%), and from 66.23 to 66.23 on the 38B model (+0.40%–2.24%). These gains remain consistently positive across all four backbones, even as baseline capabilities increase with scale. Notably, the relative improvements are more substantial at smaller model sizes (e.g., Qwen2.5-VL-7B), suggesting MindGYM is particularly effective in lower-resource or less capable models. This trend underscores the method’s architecture-agnostic and scaling-friendly design, making it highly transferable across a wide range of multimodal foundation models.
@@
Despite using only 400 synthetic training samples, MindGYM achieves performance that not only matches but exceeds baselines trained with 10$\times$ more data. This includes outperforming Open-O1 (4k) in both textual and multimodal tasks. On MathVision, MindGYM reaches 28.62 while Open-O1 (4k) collapses to 17.43. This stark contrast underscores the high quality and utility of the data synthesized by MindGYM.
Cross-modal Enhancement.
Despite containing only 400 synthetic examples in pure-text format, MindGYM significantly enhances both textual and cross-modal reasoning capabilities. As shown in Table 1, our method improves the Overall performance not only on language-only tasks, but also on multimodal benchmarks (e.g., from 62.46 to 64.95 on InternVL3-14B and 65.16 to 66.23 on InternVL3-38B). These consistent improvements are also observed in vision-intensive benchmarks like MathVista and MathVision, even though no vision data was used in fine-tuning. This demonstrates that our synthetic dataset effectively strengthens the model’s intrinsic thinking capability, enabling better generalization across modalities. Such cross-modal enhancement, achieved through text-only supervision, highlights the thinking-centered nature of our data and its ability to propagate improvements to both language and vision domains.
Summary.
These results collectively demonstrate that our method effectively tackles three key challenges in multimodal instruction tuning: data inefficiency, modality transferability, and scalability across different models and sizes. First, MindGYM achieves consistent and significant gains across all evaluation metrics with only 400 synthetic Chinese text-only samples, outperforming baselines trained with up to 10× more data, highlighting its exceptional data efficiency. Second, the improvements span both textual and vision-intensive benchmarks, even without visual supervision, indicating that our thinking-centric text data successfully enhances intrinsic thinking that transfers across modalities. Third, our method remains robust and effective across a range of backbones (Qwen2.5-VL and InternVL3) and model sizes (7B–38B), making it highly scalable and architecture-agnostic. These findings validate that our approach addresses the limitations of prior methods in terms of modality inconsistency, generalization, and computational efficiency (as outlined in Section 2).
4.3 Data analysis
Table 2: Data-Juicer analysis results across different models.
Model Dataset
quality
-mean
$\uparrow$
quality
-std
$\downarrow$
action $\uparrow$ dependency $\uparrow$ token length
baseline Openo1-sft (400) openo1 0.96 0.091 8.05 2.00 77 274
Openo1-sft (4k) openo1 0.70 0.10 8.78 2.02 83 284
LIMO (817) ye2025limo 0.87 0.10 7.71 2.02 101 322
MMEvol-SciQA (106) luo2024mmevol 0.91 0.082 2.31 1.85 13.73 67
MMEvol-DvQA (4k) luo2024mmevol 0.76 0.11 2.66 1.95 14.82 69
Qwen2.5-VL-7B MindGYM-Text (ours) 0.93 0.055 10.34 2.17 147 110
Qwen2.5-VL-32B MindGYM-Text (ours) 0.98 0.031 10.94 2.25 168 134
InternVL-8B MindGYM-Text (ours) 0.97 0.044 8.64 2.16 233 195
InternVL-38B MindGYM-Text (ours) 0.96 0.039 7.66 2.07 118 92
To comprehensively assess the quality of our MindGYM synthetic dataset, we adopt Data-Juicerdj2  [^2] https://github.com/modelscope/data-juicer, a modular data-centric analysis toolkit that provides LLM-guided operators for probing data across various dimensions. Specifically, we compare MindGYM against several baselines (e.g., LIMO, Open-O1, MMEvol) using five diagnostic filters: (1) quality, which estimates the overall textual quality based on LLM judgment; (2) action, counting the number of action verbs as a proxy for instruction richness; (3) dependency, which penalizes the presence of non-independent noun phrases based on syntactic trees; (4) token, evaluating token-length consistency; and (5) length, which considers raw text length variability. Table 2 summarizes the full results across multiple model series and checkpoints. Our key findings are elaborated below.
Table 3: The results of win rates in terms of relative improvements on MindGYM over baselines, using GPT-4 as a scorer.
Datasets Depth Breadth Avg
Raw 10.2% 19.4% 14.8% $\uparrow$
LIMO 1.63% 23.8% 12.7% $\uparrow$
Open-O1 (400) 1.88% 23.7% 12.8% $\uparrow$
MMEvol-DVQA 8.41% 37.0% 22.7% $\uparrow$
Overall 5.53% $\uparrow$ 26.0% $\uparrow$ 15.8% $\uparrow$
Higher Quality.
We observe a strong positive correlation between the average quality score (as estimated by Data-Juicerdj2 ) and downstream finetuning performance. In the baseline group, Open-O1(400) and MMEvol-SciQA achieve high quality scores (0.96 and 0.91, respectively), and correspondingly rank first and second (excluding MindGYM) in Table 1, outperforming other baselines. Notably, Open-O1(400) not only obtains the highest quality score among baselines, but also shows consistent Overall score improvements across all models compared to the raw version. MindGYM further surpass these baselines in quality: Qwen2.5-VL-7B reaches 0.93, while Qwen2.5-VL-32B achieves 0.98. These results suggest that larger LLMs not only produce better answers, but also generate higher-quality synthetic data. This trend holds across models and highlights the value of scaling for data generation.
Lower Variance.
Beyond high average quality, we observe a key phenomenon: as the parameter scale of the LLM used for data synthesis increases, the variance of quality scores decreases. For instance, Qwen2.5-VL-32B and InternVL3-38B exhibit low standard deviations (0.031 and 0.039, respectively), in contrast to baselines like MMEvol-DvQA, which shows a much higher variance (0.11) and ranks among the worst-performing in Table 1. Even within the baseline group, Open-O1(400) and MMEvol-SciQA, both with relatively lower variances, also achieve stronger Overall scores, typically outperforming their raw counterparts.
Interestingly, although MMEvol-SciQA has slightly lower variance than Open-O1(400), its downstream performance lags behind, for example, showing negative Overall gain under Qwen2.5-VL-32B. This can be attributed to its lower mean quality score, reinforcing that both high average quality and low variance are critical to data effectiveness. The low-variance patterns observed in larger models suggest their superior ability not only to generate high-quality content but also to suppress outliers and noise—an essential property for stable and robust fine-tuning. These findings underscore a key insight: consistency matters as much as mean quality in building high-impact synthetic datasets.
Richer Semantics.
We next analyze linguistic properties that reflect structural complexity and task grounding: the average number of action verbs and the count of non-independent entities derived from dependency trees. MindGYM samples show significantly higher values in both metrics. For instance, Qwen2.5-VL-32B samples contain 14.14 action verbs and 2.14 non-independent entities on average, while LIMO and MMEvol-SciQA yield only 7.71/2.02 and 2.31/1.85, respectively. Meanwhile, MindGYM maintains a compact format, averaging 133–215 tokens and 92–110 characters per sample across models. This balance of rich semantics and concise form enhances model understanding while improving supervision efficiency and signal clarity. These results suggest that MindGYM prompts are both procedurally rich and semantically dense, which facilitates stronger self-consistency in instruction following.
GPT-based Thinking Quality Scoring.
We establish a two-dimensional evaluation framework leveraging GPT-4 as an expert scorer. The protocol assesses: (1) thinking Depth via derivation complexity, and (2) thinking Breadth by solution diversity. Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix B.2.5. Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark. The experimental results documented in Table 3 demonstrate MindGYM’s consistent superiority across both evaluation axes, particularly achieving 15.8% average improvements in thinking depth over the raw model and other competitive baselines.